If 5 times the smaller of two numbers is subtracted from twice the larger, the result is 16. If the larger is increased by 3 times the smaller, the result is 63. Find the numbers.

Answer by checkley79(3341)

LET X BE THE SMALLER & Y BE THE LARGER NUMBER.
2Y-5X=16
3X+Y=63 OR Y=63-3X
2(63-3X)-5X=16
126-6X-5X=16
-11X=16-126
-11X=-110
X=-110/-11
X=10 ANS. FOR THE SMALLER NUMBER.
Y=63-3*10
Y=63-30
Y=33 ANS. FOR THE LARGER NUMBER.
PROOF:
3*10+33=63
30+33=63
63=63